6m+n=-1
2m+3n=-19

Answer by checkley77(12844)

6m+n=-1 MULTIPLY BY -3 & ADD.
2m+3n=-19
-18m-3n=3
-----------
-16m=-16
m=-16/-16
m=1 ans.
6*1+n=-1
6+n=-1
n=-1-6
n=-7 ans.
proof:
2*1+3(-7)=-19
2-21=-19
-19=-19